OpenStudy (anonymous):

Andreas received a new book for his birthday. He read three-eighths of the book in the first week, then fifty-three more pages in the second week. If he still has one hundred seventeen pages to read, how many total pages are in his book?

OpenStudy (anonymous):

Did you figure it out???

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Where are you stuck

OpenStudy (anonymous):

the 3/8ths...

OpenStudy (anonymous):

53+107=?

OpenStudy (anonymous):

I got that 117 and 53 is 170 but I don't know what to do with the 3/8

OpenStudy (anonymous):

160, therefore a number + 3/8 has to = 160, not 170

OpenStudy (anonymous):

dude 53+117=170

OpenStudy (anonymous):

You do the problem backwards. you start with the end and work your way forward...

OpenStudy (anonymous):

272

OpenStudy (anonymous):

Total number of pages in the book, we'll call that n. n=(3/8)n+53+117=...? Algebraically manipulate the function.

OpenStudy (anonymous):

How do I figure that out???

OpenStudy (anonymous):

I tried but I just can't get it

OpenStudy (anonymous):

Alright, so let's start with n=(3/8)n+53+117. n is the total number of pages. We read 3/8 of n, so (3/8)n. Then we read 53 more pages. Then we still have 117 pages more to read. A total of n. Yes? Subtract (3/8)n from both sides. (5/8)n=53+117. Add the right side. 170. Divide by 5. 34. Multiply by 8. You get n=...?

OpenStudy (anonymous):

I am not sure that is correct

OpenStudy (anonymous):

I thought that you subtracted 53 from both sides and then divide 3/8 from both sides... Am i correct?